OpenStudy (anonymous):

what is the thermal capacity of 3kg of water?

OpenStudy (anonymous):

thermal capacity = mass of substance * specific heat capacity of the substance =m*s For water specific heat,s=4186 J/kg/K and we are given m=3 kg

OpenStudy (anonymous):

isnt thermal capacity same as specific heat capacity?

OpenStudy (anonymous):

@mr.singh thermal capacity (also known as heat capacity) is different from specific heat capacity. Heat capacity is an extensive property i.e. depends upon the mass of the substance you have taken while specific heat is an intensive property of the type of material you have taken. The above formula is a relation between them!